With the aid of the current, Joe can row a canoe 3 miles in 12 minutes. Against the current he requires 18 minutes to row the same distance. How fast does Joe row in still water and how fast is the current?
Answer by Alan3354(69443)

With the aid of the current, Joe can row a canoe 3 miles in 12 minutes. Against the current he requires 18 minutes to row the same distance. How fast does Joe row in still water and how fast is the current?
--------------
3 miles in 12 minutes = 15 mph
3 miles in 18 minutes = 10 mph
His rowing speed is the average = (15+10)/2 = 12.5 mph
The current speed is the difference = 15 - 12.5 = 2.5 mph